OpenStudy (anonymous):

Jackson put twelve gallons of gasoline in his dad”s car and spent twenty-four dollars. Use a direct variation to determine how many gallons of gasoline he put in his mom's car if he spent twenty dollars.

OpenStudy (ragingsquirrel):

10 gallons for $20 i believe.

OpenStudy (anonymous):

12/24 = 1/2 = ?/20

OpenStudy (anonymous):

Formula time Total Cost = gallons * cost per gallon